# On the kernels of the pro-$p$ outer Galois representations associated to once-punctured CM elliptic curves

### v2

@@ 4.2. Group-theoretic lemmas
In this appendix, we keep the same notation as in the main text, i.e. let $K$ be an imaginary quadratic field of class number one, $p$ a prime $\geq 5$ which splits in $K$ as $(p)=\mathfrak{p}\bar{\mathfrak{p}}$ and $\mu_{p}$ the group of $p$-th roots of unity. Moreover, let $K_{\infty}$ be the unique $\mathbb{Z}_{p}^{2}$-extension of $K$. In this appendix, we determine the structure of the Galois group of the maximal pro-$p$ extension $\Omega^{\mathrm{cyc}}_{K}$ of $K(\mu_{p})$ unramified outside $p$ when $p$ satisfies a certain assumption.
@@
We denote a unique prime of $K(\mu_{p})$ lying above $\mathfrak{p}$ (resp. $\bar{\mathfrak{p}}$) by the same letter, and fix an arbitrary prime of $\Omega_{K}^{\mathrm{cyc}}$ above $\mathfrak{p}$. Then we have the homomorphism
$\phi_{\mathfrak{p}}:G_{\mathbb{Q}_{p}(\mu_{p})}^{(p)}\to\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p})),$
associated to the decomposition group of the fixed prime of $\Omega_{K}^{\mathrm{cyc}}$.
@@
The Galois group $\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))$ is said to be local type if $\phi_{\mathfrak{p}}$ is surjective [NSW08, (10.9.6) Definition]. Moreover, it is called purely local if $\phi_{\mathfrak{p}}$ is an isomorphism [Win90]. The aim of this appendix is to prove the following theorem, which tells us exactly when $\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))$ is purely local:
@@
The Galois group $\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))$ is purely local if and only if $p$ does not divide the class number of $K(\mu_{p})$ and there exists a unique prime of $K_{\infty}$ lying above $\mathfrak{p}$.
Remark A.2.
(1) The Galois group $G_{\mathbb{Q}(\mu_{p})}^{(p)}$ is a pro-$p$ Demushkin group of rank $p+1$ [NSW08, (7.5.11) Theorem (ii)]. In particular, it is generated by $p+1$ elements which are subject to a single relation. On the other hand, if $p$ does not divide the class number of $K(\mu_{p})$, then $\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))$ is also generated by $p+1$ elements satisfying a single relation [NSW08, (10.7.13) Theorem]. Hence the condition on the class number implies that the concerned Galois groups have the same numbers of generators and relations.
(2) The latter condition on the number of primes of $K_{\infty}$ above $\mathfrak{p}$ fails when $K=\mathbb{Q}(\sqrt{-1})$ and $p=29789$ [Ish25, Example 4.16]. This is the only example the author has been able to find where the latter condition does not hold when $K=\mathbb{Q}(\sqrt{-1})$.
@@
Assume that $\phi_{\mathfrak{p}}$ is an isomorphism. Then there exists a unique prime of $\Omega^{\mathrm{cyc}}_{K}$ above $\mathfrak{p}$. Since we have $K_{\infty}\subset K_{\infty}K(\mu_{p})\subset\Omega^{\mathrm{cyc}}_{K}$, a prime of $K_{\infty}$ lying above $\mathfrak{p}$ is also unique. Moreover, the class number of $K(\mu_{p})$ is not divisible by $p$ by [Ish25, Proposition 4.13].
@@
$H^{1}(\mathrm{Gal}(\Omega^{\mathrm{cyc}}_{K}/K(\mu_{p})),\mu_{p})\to H^{1}(G_{\mathbb{Q}_{p}(\mu_{p})}^{(p)},\mu_{p})\xrightarrow{\sim}\mathbb{Q}_{p}(\mu_{p})^{\times}/p,$
induced by $\phi_{\mathfrak{p}}$. Since the class number of $K(\mu_{p})$ is prime to $p$, we have
$O_{K(\mu_{p})}[1/p]^{\times}/p\xrightarrow{\sim}H^{1}(\mathrm{Gal}(\Omega^{\mathrm{cyc}}_{K}/K(\mu_{p})),\mu_{p})$
by using a similar argument as the proof of Lemma 4.2. Hence the surjectivity of $\phi_{\mathfrak{p}}$ is equivalent to the injectivity of the map
@@
induced by the $\mathfrak{p}$-adic completion. The map $\iota_{\mathfrak{p}}$ is compatible with the action of $\mathrm{Gal}(K(\mu_{p})/K)$, the decomposition group of $\mathfrak{p}$. Moreover, if we define
@@
to be the map induced by the $\bar{\mathfrak{p}}$-adic completion, then the map
@@
is injective, since the dimension of the kernel of $(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}})$ is equal to that of the $p$-class group of $K(\mu_{p})$ modulo $p$ (see [NSW08, Proof of (10.7.3)], for example).
The injectivity of $(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}})$ already implies that the kernel of $\iota_{\mathfrak{p}}$ is at most a single Tate twist $\mathbb{F}_{p}(1)$: More precisely, we show that the composite
@@
is injective, where $K(\mu_{p})^{+}$ denotes the maximal real subfield of $K(\mu_{p})$. Note that
$\dim_{\mathbb{F}_{p}}O_{K(\mu_{p})^{+}}[1/p]^{\times}/p=p\quad\text{and}\quad\dim_{\mathbb{F}_{p}}O_{K(\mu_{p})}[1/p]^{\times}/p=p+1$
by Dirichlet’s unit theorem. Moreover, the cokernel of the inclusion
$O_{K(\mu_{p})^{+}}[1/p]^{\times}/p\to O_{K(\mu_{p})}[1/p]^{\times}/p$
is generated by the image of $\mu_{p}$. The injectivity of $\iota_{\mathfrak{p}^{+}}$ follows from the that of the map $(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}})$ and the fact that a prime of $K(\mu_{p})^{+}$ lying above $p$ is unique, since the restriction of $(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}})$ on $O_{K(\mu_{p})^{+}}[1/p]^{\times}/p$ is just the diagonal map which is induced by the completion at the unique prime $K(\mu_{p})^{+}$ lying above $p$.
@@
as an $\mathbb{F}_{p}[\mathrm{Gal}(K(\mu_{p})/K)]$-module. Since $\ker(\iota_{\mathfrak{p}})$ is stable under the action of $\mathrm{Gal}(K(\mu_{p})/K)$, the injectivity of $\iota_{\mathfrak{p}^{+}}$ implies that $\ker(\iota_{\mathfrak{p}})\subset\mathbb{F}_{p}(1)^{\oplus 2}$. In terms of Galois groups, the cokernel of the map
$G_{\mathbb{Q}_{p}(\mu_{p})}^{(p),\mathrm{ab}}/p\to\mathrm{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))^{\mathrm{ab}}/p$
induced by $\phi_{\mathfrak{p}}$ is contained in $\mathbb{F}_{p}^{2}$ (with trivial $\mathrm{Gal}(K(\mu_{p})/K)$-action). However, since the $\mathbb{F}_{p}^{2}$-extension of $K(\mu_{p})$ corresponding to $\mathbb{F}_{p}(1)^{\oplus 2}$ is given by the quotient
$\mathrm{Gal}(K_{\infty}K(\mu_{p})/K(\mu_{p}))\cong\mathbb{Z}_{p}^{2}\to\mathbb{F}_{p}^{2}$
and we have assumed that there is a unique prime of $K_{\infty}K(\mu_{p})$ lying above $\mathfrak{p}$, the decomposition group at $\mathfrak{p}$ must coincide with $\mathbb{F}_{p}^{2}$. Hence $\phi_{\mathfrak{p}}$ is surjective as desired. ∎
Theorem A.1 allows us to use the local Tate duality to compute cohomology groups of the global Galois group $\mathrm{Gal}(\Omega^{\mathrm{cyc}}_{K}/K(\mu_{p}))$. Consequently, we obtain the following corollary:
@@
Under the same assumption as in Theorem A.1, the cohomology group $H^{2}_{\mathrm{\acute{e}t}}(O_{K}[1/p],\mathbb{Z}_{p}(m_{1},m_{2}))$ is finite for every $(m_{1},m_{2})\in I$ such that $m_{1}\equiv m_{2}\bmod p-1$. Consequently, the $(m_{1},m_{2})$-th elliptic Soulé character $\kappa_{(m_{1},m_{2})}$ is nontrivial for every such $(m_{1},m_{2})\in I$.

### v3

@@ Appendix A Pure locality of $p$-th cyclotomic extension of imaginary quadratic field
We keep the same notation as in the main text: Let $K$ be an imaginary quadratic field of class number one, $p$ a prime $\geq 5$ which splits in $K$ as $(p)=\mathfrak{p}\bar{\mathfrak{p}}$, $\mu_{p}$ the group of $p$-th roots of unity and $K_{\infty}$ the $\mathbb{Z}_{p}^{2}$-extension of $K$. In this appendix, we determine the structure of the Galois group of the maximal pro-$p$ extension $\Omega^{\mathrm{cyc}}_{K}$ of $K(\mu_{p})$ unramified outside $p$ under a certain assumption.
@@
We denote a unique prime of $K(\mu_{p})$ lying above $\mathfrak{p}$ (resp. $\bar{\mathfrak{p}}$) by the same letter, and fix an arbitrary prime of $\Omega_{K}^{\mathrm{cyc}}$ lying above $\mathfrak{p}$. We obtain a homomorphism
$\phi_{\mathfrak{p}}:G_{\mathbb{Q}_{p}(\mu_{p})}^{(p)}\to\operatorname{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p})),$
@@
associated to the decomposition group of the fixed prime of $\Omega_{K}^{\mathrm{cyc}}$. The Galois group is said to be purely local if $\phi_{\mathfrak{p}}$ is an isomorphism [Win90] (see also [NSW08, (10.9.6) Definition]). The aim of the present appendix is to prove the following, which tells us exactly when the Galois group is purely local:
@@
The Galois group $\operatorname{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))$ is purely local if and only if $p$ does not divide the class number of $K(\mu_{p})$ and there exists a unique prime of $K_{\infty}$ above $\mathfrak{p}$.
Remark A.1.1.
(1) The Galois group $G_{\mathbb{Q}(\mu_{p})}^{(p)}$ is a pro-$p$ Demushkin group of rank $p+1$ [NSW08, (7.5.11) Theorem (ii)]. In particular, it is generated by $p+1$ elements that satisfy one relation. On the other hand, if $p$ does not divide the class number of $K(\mu_{p})$, then $\operatorname{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))$ is also generated by $p+1$ elements that satisfy one relation [NSW08, (10.7.13) Theorem]. Hence the first condition on the class number implies that both Galois groups have the same numbers of generators and relations.
(2) The latter condition on the number of primes of $K_{\infty}$ above $\mathfrak{p}$ fails when $K=\mathbb{Q}(\sqrt{-1})$ and $p=29789$ [Ish25, Example 4.16]. This is the only example the author has been able to find where the latter condition does not hold for $K=\mathbb{Q}(\sqrt{-1})$.
@@
Assume that $\phi_{\mathfrak{p}}$ is an isomorphism. Then there exists a unique prime of $\Omega^{\mathrm{cyc}}_{K}$ lying above $\mathfrak{p}$. Since $K_{\infty}$ is a subfield of $\Omega^{\mathrm{cyc}}_{K}$, a prime of $K_{\infty}$ lying above $\mathfrak{p}$ is uniquely determined. Moreover, the class number of $K(\mu_{p})$ is not divisible by $p$ by [Ish25, Proposition 4.13].
@@
$H^{1}(\operatorname{Gal}(\Omega^{\mathrm{cyc}}_{K}/K(\mu_{p})),\mu_{p})\to H^{1}(G_{\mathbb{Q}_{p}(\mu_{p})}^{(p)},\mu_{p})\xrightarrow{\sim}\mathbb{Q}_{p}(\mu_{p})^{\times}/p,$
induced by $\phi_{\mathfrak{p}}$, where the latter isomorphism comes from Kummer theory. Since the class number of $K(\mu_{p})$ is prime to $p$, Kummer theory gives an isomorphism
$O_{K(\mu_{p})}[1/p]^{\times}/p\xrightarrow{\sim}H^{1}(\operatorname{Gal}(\Omega^{\mathrm{cyc}}_{K}/K(\mu_{p})),\mu_{p}).$
Hence the surjectivity of $\phi_{\mathfrak{p}}$ is equivalent to the injectivity of the map
@@
induced by the $\mathfrak{p}$-adic completion. The map $\iota_{\mathfrak{p}}$ is compatible with the action ofthe decomposition group of $\operatorname{Gal}(K(\mu_{p})/K)$ at $\mathfrak{p}$, which is nothing but $\operatorname{Gal}(K(\mu_{p})/K)$ itself. Let
@@
be the map induced by the $\bar{\mathfrak{p}}$-adic completion. Then the map
@@
is injective, since the dimension of the kernel of $(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}})$ coincides with that of the $p$-class group of $K(\mu_{p})$ modulo $p$ (see [NSW08, Proof of (10.7.3)], for example). On the other hand, this injectivity already implies that the kernel of $\iota_{\mathfrak{p}}$ is at most a single Tate twist $\mathbb{F}_{p}(1)$: More precisely, we claim that the composite
@@
is injective, where $K(\mu_{p})^{+}$ denotes the maximal real subfield of $K(\mu_{p})$. Note that, by Dirichlet’s unit theorem, we have
$\dim_{\mathbb{F}_{p}}O_{K(\mu_{p})^{+}}[1/p]^{\times}/p=p\quad\text{and}\quad\dim_{\mathbb{F}_{p}}O_{K(\mu_{p})}[1/p]^{\times}/p=p+1,$
and the cokernel of the inclusion $O_{K(\mu_{p})^{+}}[1/p]^{\times}/p\to O_{K(\mu_{p})}[1/p]^{\times}/p$ is generated by $\mu_{p}\cong\mathbb{F}_{p}(1)$. The injectivity of $\iota_{\mathfrak{p}^{+}}$ follows from the that of the map $(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}})$ and the fact that a prime of $K(\mu_{p})^{+}$ lying above $p$ is unique, since the restriction of $(\iota_{\mathfrak{p}},\iota_{\bar{\mathfrak{p}}})$ on $O_{K(\mu_{p})^{+}}[1/p]^{\times}/p$ is regarded as the diagonal map induced by the completion at the unique prime $K(\mu_{p})^{+}$ lying above $p$.
@@
as an $\mathbb{F}_{p}[\operatorname{Gal}(K(\mu_{p})/K)]$-module. Since $\operatorname{Ker}(\iota_{\mathfrak{p}})$ is stable under the action of $\operatorname{Gal}(K(\mu_{p})/K)$, the injectivity of $\iota_{\mathfrak{p}^{+}}$ implies that $\operatorname{Ker}(\iota_{\mathfrak{p}})$ is contained in $\mathbb{F}_{p}(1)^{\oplus 2}$. In terms of Galois groups, we have proved that the cokernel of
$G_{\mathbb{Q}_{p}(\mu_{p})}^{(p),\mathrm{ab}}/p\to\operatorname{Gal}(\Omega_{K}^{\mathrm{cyc}}/K(\mu_{p}))^{\mathrm{ab}}/p$
induced by $\phi_{\mathfrak{p}}$ is contained in $\mathbb{F}_{p}^{2}$ (with trivial $\operatorname{Gal}(K(\mu_{p})/K)$-action). However, since this $\mathbb{F}_{p}^{2}$-extension is given by the modulo-$p$ quotient
$\operatorname{Gal}(K_{\infty}K(\mu_{p})/K(\mu_{p}))\cong\mathbb{Z}_{p}^{2}\to\mathbb{F}_{p}^{2}$
and we have assumed that there is a unique prime of $K_{\infty}K(\mu_{p})$ lying above $\mathfrak{p}$, the map $\phi_{\mathfrak{p}}$ is surjective. ∎
Theorem A.1 allows us to use the local Tate duality to compute cohomology groups of $\operatorname{Gal}(\Omega^{\mathrm{cyc}}_{K}/K(\mu_{p}))$. Consequently, we obtain the following corollary:
@@
We keep the same assumption as in Theorem A.1. Then the cohomology group $H^{2}_{\mathrm{\acute{e}t}}(O_{K}[1/p],\mathbb{Z}_{p}(m_{1},m_{2}))$ is finite for every $(m_{1},m_{2})\in I$ such that $m_{1}\equiv m_{2}\bmod p-1$. Consequently, the $(m_{1},m_{2})$-th elliptic Soulé character $\kappa_{(m_{1},m_{2})}$ is nontrivial for every such $(m_{1},m_{2})\in I$.